# Amputation-imputation based generation of synthetic tabular data for ratemaking

### v1

@@ 5 Conclusion
Our paper demonstrates that the amputation-imputation framework based on MICE can be effectively used within an actuarial context. This framework is competitive for generating synthetic ratemaking data when compared to more sophisticated deep generative models such as CTGANs or VAEs.
In our case studies based on the open-source freMTPL2freq dataset, synthetic data generated using MICE with random forests as the imputation engine shows, on average, the best properties in terms of modeling efficacy and achieves high ranks in model accuracy and coefficient difference metrics. In general, several of the $9$ tested methods for synthetic data generation seem to have their niches for specific use cases.
Having subjectively evaluated the ease of use for each tested method, we identify MICE as the most user-friendly approach, since it requires the least amount of additional work to get started, which is an important practical aspect for actuaries. Custom deep generative models like WGAN by Cote et al., (2020) or VAE by Jamotton and Hainaut, (2024) seem to require significantly more expertise and tuning to get good results.
Our case study also shows that generic data augmentation does not generally improve the performance of GLMs trained on the real data augmented with the synthetic one. We observe that an increase in the proportion of synthetic data compared to the original training data leads to an approximately linear increase in the metrics that measure the difference between coefficients of two GLMs, e.g., the true GLM vs a GLM estimated on augmented data. We have noticed only one case where data augmentation improved the estimation of the true GLM coefficients.
We leave several aspects for future research. For example, it may be interesting to investigate how to integrate business constraints into the amputation-imputation framework based on MICE, e.g., linear constraints on numerical variables or hierarchy among categorical variables. Another relevant topic for further research may be the privacy of synthetic datasets generated with the help of MICE.
@@ Acknowledgments
Yevhen Havrylenko acknowledges the financial support by the PRIME program (link) of the German Academic Exchange Service (DAAD, https://ror.org/039djdh30) with funds from the German Federal Ministry of Education and Research (BMBF). Yevhen Havrylenko acknowledges the support of the University of Copenhagen and Johannes Kepler University Linz, where he conducted parts of this research project. Meelis Käärik and Artur Tuttar acknowledge the support of the Estonian Research Council grant PRG1197.
The authors of this publication express their gratitude to Raul Kangro and Märt Möls for their insights, ideas, and discussion related to the article. The authors are grateful to the participants of the 28th International Congress on Insurance: Mathematics and Economics.
@@ Appendix A Appendix A
    use Nearest Centroid Classification to identify for this vector its the closest class; the centroid of each class is determined by the encoding (given by the respective AE fitted before training $GM_{A}$) of the respective category of the original categorical variable;

### v2

@@ 5 Conclusion
Our paper demonstrates that the amputation-imputation framework based on MICE can be effectively used within an actuarial context. This framework is competitive with more sophisticated deep generative models, such as CTGANs or VAEs, for generating ratemaking data.
In our case studies based on the open-source freMTPL2freq dataset, synthetic data generated using MICE with random forests as the imputation engine shows, on average, the best properties in terms of modeling efficacy and achieves high ranks in model accuracy and coefficient difference metrics. In general, several of the $10$ tested methods for synthetic data generation seem to have their niches for specific use cases.
Having subjectively evaluated the ease of use for each tested method, we identify MICE as the most user-friendly approach, since it requires the least amount of additional work to get started, which is an important practical aspect for actuaries. Custom deep generative models like WGAN by Côté et al. (2025) or VAE by Jamotton and Hainaut (2024) seem to require significantly more expertise and tuning to get good results.
Our case study also shows that generic data augmentation does not generally improve the performance of GLMs trained on the real data augmented with the synthetic one. We observe that increasing the proportion of synthetic data relative to the original training data results in approximately linear increases in metrics measuring the difference between the coefficients of two GLMs, e.g., the true GLM vs a GLM estimated on augmented data. We have observed only one case in which data augmentation improved the estimation of the true GLM coefficients.
We leave several aspects for future research. For example, it may be interesting to investigate how to integrate business constraints into the amputation-imputation framework based on MICE, e.g., linear constraints on numerical variables or hierarchy among categorical variables. Another research direction is the exploration of data augmentation when the original training dataset is small or when the synthetic data generator contains information not present in the training data. While ratemaking has historically relied on association measures like correlation, future research must rigorously evaluate the extent to which synthetic data generators preserve the underlying causal structure of insurance risk. Finally, future studies should systematically assess the disclosure risk inherent in each generator, specifically quantifying their vulnerability to membership inference and reconstruction attacks.
@@ Acknowledgments
Yevhen Havrylenko acknowledges the financial support by the PRIME program (link) of the German Academic Exchange Service (DAAD, https://ror.org/039djdh30) with funds from the German Federal Ministry of Education and Research (BMBF). Yevhen Havrylenko acknowledges the support of Ulm University, the University of Copenhagen and Johannes Kepler University Linz, where he conducted parts of this research project. Meelis Käärik and Artur Tuttar acknowledge the support of the Estonian Research Council grants PRG1197 and PRG3105.
The authors of this publication express their gratitude to Raul Kangro and Märt Möls for their insights, ideas, and discussion related to the article. The authors are grateful to the participants of the 28th International Congress on Insurance: Mathematics and Economics and 1st ASTIN Bulletin conference. The authors are thankful to Magnus Käärik for assistance in refining the language and enhancing clarity.
@@
Data availability statement
The authors intend to upload their code and data to a publicly accessible GitHub repository upon acceptance of the article for publication.
@@ Appendix A Appendix A
    use Nearest Centroid Classification to identify for this vector its closest class; the centroid of each class is determined by the encoding (given by the respective AE fitted before training $GM_{A}$) of the respective category of the original categorical variable;